OpenStudy (anonymous):

Laplace transform of a function f(x) is function (Lf)(s) of the variable s defined by the improper integral (Lf)(s)= (definite integral [0,infinity)) f(x)e^(-sx) dx "s" is a constant and let "a" be a real number. Find: 1.) the sine function f(x)=sin(ax) 2.)The exponential function f(x)=e^(ax). assume s>a 3.) the cosine function f(x)=cos(ax). assume s>a

OpenStudy (anonymous):

Integration by parts for all of these integrals.

zepdrix (zepdrix):

Which one are you stuck on? All of them? :o

OpenStudy (anonymous):

well, I never did them before so if i can see one I can try to teach from the notes. I did a basic one with a constant time e^-sx and did that fine but these are a little more in depth.

OpenStudy (anonymous):

*constant times e^-sx

zepdrix (zepdrix):

Were you able to get through the constant * e^-sx without too much trouble? Because part b will work out almost exactly the same way, but with a minor change.

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\LARGE \mathscr{L}\left\{e^{ax}\right\}\quad=\quad \int\limits_0^{\infty}e^{ax}\cdot e^{-sx}dx\]

zepdrix (zepdrix):

Rules of exponents lets us write it like this:\[\LARGE \mathscr{L}\left\{e^{ax}\right\}\quad=\quad \int\limits\limits_0^{\infty}e^{-sx+ax}dx\]We'll further simplify the exponent to:\[\LARGE =\quad \int\limits\limits\limits_0^{\infty}e^{-(s-a)x}\;dx\]

zepdrix (zepdrix):

So when you evaluated,\[\LARGE =\quad \int\limits\limits\limits\limits_0^{\infty}e^{\color{orangered}{-s}x}\;dx\]What did you come up with? :)

OpenStudy (anonymous):

-(e^(-sx))/s

OpenStudy (anonymous):

bound [0,ininity)

zepdrix (zepdrix):

Ok good, and evaluating it at the limits should simplify it down really nicely.

OpenStudy (anonymous):

pulled out 1/-s cause constant

OpenStudy (anonymous):

for the others i would use IPB?

zepdrix (zepdrix):

Others? :o Wait wait, I'm confused, what did you actually end up with for the laplace of 1

zepdrix (zepdrix):

You have to "plug in" the limits and it should simplify down. It's not -1/s, but that's close D:

OpenStudy (anonymous):

oh, i know that. The one problem that i finished before is part 4, I did not list it because I was able to solve it. The problem you just helped me with, i did not think of the exponent law and was trying to IPB. I wanted to know if i would use IPB for f(x)=sin(ax) and fx=cosax.

OpenStudy (anonymous):

would my u be e^(-sx) and dv be sin(ax) dx?

zepdrix (zepdrix):

It'll be a little tricky, we'll have to do integration by parts `two times` and then do some fancy algebra. It doesn't matter which you choose to be `u` and `dv`, we'll just have to make a note of it for the next time we do our parts, because we have to do them in the same manner we did the first time. If that doesn't make sense, that's ok. :) Yes, let u=the exponential and dv=the trig function.

zepdrix (zepdrix):

Are you familiar with Euler's Formula? There is a way to do this that is a bit easier using the complex exponential. We would convert our sine or cosine to exponentials using Euler and then the problem would work out similar to part 2. But if you haven't learned about that stuff yet, np. :)

OpenStudy (anonymous):

sorry, I never learnt it yet

zepdrix (zepdrix):

So we're working on the sine problem right? After integrating once, I think we end up with this.\[\Large -\frac{1}{s}\sin(ax)e^{-sx}-\frac{a}{s}\int\limits \cos(ax)e^{-sx}\;dx\] If I'm moving too fast, you can let me know. I don't wanna take the fun away if you're doing this on paper, hehe.

OpenStudy (anonymous):

yeah, I'm on paper cause I have to write this problem up in a 8 page explanation on how it is done.

OpenStudy (anonymous):

na, it's a good speed

zepdrix (zepdrix):

Don't worry about the limits for now, we'll take care of the whole mess at the end.

zepdrix (zepdrix):

So ummm...

zepdrix (zepdrix):

Ahh crap I did u=trig function, dv=exponential :( So yours will end up looking different. My bad, lemme fix that.

OpenStudy (anonymous):

i was just going to ask about that

OpenStudy (anonymous):

ha

zepdrix (zepdrix):

So if you do it the way you had suggested, I think we come up with:\[\Large =-\frac{1}{a} \cos(ax)e^{-sx}-\frac{s}{a}\int\limits\limits \cos(ax)e^{-sx}\;dx\]

OpenStudy (anonymous):

yes i got that

zepdrix (zepdrix):

Ok this is the part where we need to be careful. We're going to do IBP again on the integral that is remaining. Remember how we chose to let, `u=exponential` , `dv=trig function` ? We need to that here again. If we chose them in the reverse order, we would simply undo our last IBP, which is no good :p

OpenStudy (anonymous):

got it

zepdrix (zepdrix):

\[\Large =-\frac{1}{a} \cos(ax)e^{-sx}\color{orangered}{-\frac{s}{a}\int\limits\limits\limits \cos(ax)e^{-sx}\;dx}\]So I guess we'll end up with:\[\large =-\frac{1}{a} \cos(ax)e^{-sx}\color{orangered}{-\frac{s}{a}\left[\frac{1}{a}\sin(ax)e^{-sx}+\frac{s}{a}\int\limits \sin(ax)e^{-sx}\right]}\]

OpenStudy (anonymous):

good. is this one of the problem were we set the original integral equal to a variable and since integral that is in red bottom is the same we sub in that variable and solve?

zepdrix (zepdrix):

Yes, it's one of those things :)

OpenStudy (anonymous):

like w= bla bla bla +s/a w

OpenStudy (anonymous):

nice

zepdrix (zepdrix):

exactly. But don't forget to `distribute` that -s/a in front of the square brackets. So we really have -(s^2/a^2) W on the right there.

OpenStudy (anonymous):

okay

zepdrix (zepdrix):

So as you mentioned, we end up with:\[\Large W\quad=\quad=-\frac{1}{a} \cos(ax)e^{-sx}-\frac{s}{a^2}\sin(ax)e^{-sx}-\frac{s^2}{a^2}W\]

zepdrix (zepdrix):

\[\Large W+\frac{s^2}{a^2}W\quad=\quad W\left(1+\frac{s^2}{a^2}\right)\]Then we'll just divide by that coefficient on W, right? :)

OpenStudy (anonymous):

that big mess 2 post above with out the end term of W divided by (1+s^2/a^2).

zepdrix (zepdrix):

Oh I guess getting a common denominator would be good :P heh\[\Large W\left(\frac{s^2+a^2}{a^2}\right)\]

OpenStudy (anonymous):

what is the purpose of the common denominator?

zepdrix (zepdrix):

It makes it MUCH easier to divide it over to the other side. Dividing the fraction is the same as multiplying the other side by:\(\Large \left(\cfrac{a^2}{s^2+a^2}\right)\)

OpenStudy (anonymous):

okay

zepdrix (zepdrix):

See how the a's on cosine and sine will cancel out nicely when we do that? :o

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\Large W\quad=\quad -\frac{a}{s^2+a^2}\cos(ax)e^{-sx}-\frac{1}{s^2+a^2}\sin(ax)e^{-sx}\]

zepdrix (zepdrix):

Mmmmm k, giving us something like that.

zepdrix (zepdrix):

Make sure you understand up to that point. Our next step is to FINALLY "plug in" our limits of integration.

OpenStudy (anonymous):

I fine up to now

OpenStudy (anonymous):

b-a

zepdrix (zepdrix):

It looks like the upper limit of integration will be producing 0 for both terms because:\[\LARGE \lim_{x\to\infty}e^{-sx}\quad=\quad 0\] And we have that exponential appearing in both terms.

zepdrix (zepdrix):

Hmm you might want to be a little bit more rigorous though in that part. Like maybe mention something about the fact that sine and cosine are bounded by -1 and 1 so we don't need to worry about an indeterminate form. I dunno, maybe something like that D:

zepdrix (zepdrix):

We `subtract` the function evaluated at the lower limit. See how each term already has a negative in front? So we're subtracting each of those negative terms, so they'll change to positive, leaving us with:\[\Large \frac{a}{s^2+a^2}\cos(0)e^{0}+\frac{1}{s^2+a^2}\sin(0)e^{0}\]

OpenStudy (anonymous):

so a/s^2 + a^2

zepdrix (zepdrix):

Yayyy good job team! \c:/

OpenStudy (anonymous):

thanks. And how would the (assume s>a) come into play for the #2 when you showed me?

zepdrix (zepdrix):

It applies the same way it does to integrating constants, so make sure you understand that little detail.\[\Large \mathscr{L}\{1\}\quad=\quad \int\limits\limits\limits_0^{\infty}e^{-sx}\;dx\quad=\quad -\frac{1}{s}e^{-sx}|_0^{\infty}\]\[\Large =\quad -\frac{1}{s}\left[\color{red}{\lim_{b\to\infty}e^{-sb}}-e^{0}\right]\]

zepdrix (zepdrix):

Oh so for the constant case, we say that s > 0. If s is allowed to be negative, then this red part, our upper limit, would not converge to 0. Because we would have e^{+infinity} which blows up.

zepdrix (zepdrix):

The same thing happens with the s>a problem. We need the coefficient in front of b to remain negative. If it switches to positive we have a problem.

OpenStudy (anonymous):

okay. And for the f(x)= cosax it is basically what we just did right just working with cos

zepdrix (zepdrix):

Yah it'll be the exact same process. We'll just end up different coefficients in front of each term along the way.

OpenStudy (anonymous):

okay. Thanks for all of the help. You really saved me with this.

zepdrix (zepdrix):

np \c:/

OpenStudy (anonymous):

oops ha repost

OpenStudy (anonymous):

thanks again and good night. I have a ton of writing to do ha